To: J. Albertine, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IR FEL Weekly Report: August 26-30, 1996
Date: August 30, 1996
- A detailed agenda was assembled for the LPC workshop to be
held on Sept. 9-10th at Jefferson Lab. Following status reports
on the Consortium and the FEL construction, a series of talks
will be presented on laser-materials interactions in the IR.
The working groups established in the January LPC meeting will
continue their planning activities for User Lab fit-up and experiments.
Injector Test Stand
- Commissioning of the 500kV beam line for the Injector Test
Stand continued. Using beam extracted from the GaAs photocathode
with a HeNe drive laser, the transverse emittance measurement
system was tested and the design dispersion for the magnetic
spectrometer was verified. Switching to the doubled Nd:YLF (green)
drive laser, the system successfully delivered a 22 microamp beam
to the beam dump with the drive laser operated at a total power
of 5mW at a 75MHz repetition rate. With the drive laser power
increased to 0.65 watts, a 1mA average current beam was delivered
to the dump. The peak charge per bunch was in the range of 25pC.
With this successful test of all the beamline components and
instrumentation, the gun assembly will be taken off line next
week to change out the GaAs cathode. The present cathode has
a crack and is misaligned but we were able to work around these
problems by manipulation of the drive laser alignment.
- The first coupons of ion implanted alumina from LBL were tested
as an alternative coating method for the new ceramic stack assembly
for the ITS. New samples with higher implantation doses are expected
next week. The new cathode support tube was completed and is
ready for polishing.
- Work continues on design of the 180 degree dipole at Northrop
Grumman with completion expected to be per schedule. The prototype
major dipole was assembled and the magnetic measurements test
stand prepared to start tests. The design of the weak quadrupole
was qualified by magnetic testing.
- The standard CEBAF klystrons were operated at 8kW average
power for two days with no problems arising. The procurement
cycle for a High Power Amplifier (HPA) that will allow operation
of the IRFEL cryomodule at this increased power level was initiated.
- Drawings for the optical cavity components are in internal
review and check-off. A statement of work was drafted for CNU
for the design of an interferometric cavity length measurement
- The scheduled date for pouring the final third of the FEL
Building floor has been rescheduled to Sept. 2. Several non-conformances
on placement of support columns and the pass-through culverts
for the cryogenic lines had to be resolved prior to the final
pour and start of the wall sections.